# The Impact of the Hubble Tension on the Evidence for Dynamical Dark Energy

### v1

@@ Abstract
Recent findings from the Dark Energy Spectroscopic Instrument (DESI) Data Release 2 (DR2) favor a dynamical dark energy characterized by a phantom crossing feature. This result also implies a lower value of the Hubble constant, thereby intensifying the so-called Hubble tension. To alleviate the Hubble tension, we consider the early dark energy and explore its impact on the evidence for dynamical dark energy including the Hubble constant calibrated by the SH0ES collaboration. We find that incorporating SH0ES prior with CMB, DESI DR2 BAO and Pantheon Plus/Union3/DESY5 data reduces the preference to dynamical dark energy to $1.8\sigma/1.8\sigma/2.9\sigma$ level, respectively. Our results suggest a potential tension between the Hubble constant $H_{0}$ of the SH0ES measurement and the phantom-to-quintessence transition in dark energy favored by DESI DR2 BAO data.
@@ I Introduction
Furthermore, the constrains on the $w_{0}w_{a}$ model indicate an evolutionary behavior in the dark energy equation of state, transitioning from $w<-1$ (phantom-like Caldwell (2002)) at high redshifts to $w>-1$ (quintessence-like Ratra and Peebles (1988); Ferreira and Joyce (1998); Wetterich (1988); Scherrer and Sen (2008)) at low redshifts. This behavior implies that the dark energy density decreases with the expansion of the Universe at low redshifts, ultimately resulting in a lower value for the Hubble constant $H_{0}$. As reported in Abdul Karim et al. (2025b), the Hubble constant is determined to be $H_{0}=63.6^{+1.6}_{-2.1}\;\mathrm{km\ s^{-1}Mpc^{-1}}$ in the $w_{0}w_{a}$ model, constrained by CMB+DESI DR2 BAO. This value exhibits a tension of $4.9\sigma$ with the SH0ES measurement ($H_{0}=73.04\pm 1.04$ $\mathrm{km\ s^{-1}Mpc^{-1}}$) Riess et al. (2022). Furthermore, when considering data combinations such as CMB+DESI DR2 BAO+Pantheon Plus/Union3/DESY5, the corresponding constraints are found to be $H_{0}=67.51\pm 0.59$, $65.91\pm 0.84$, and $66.74\pm 0.56\;\mathrm{km\ s^{-1}Mpc^{-1}}$. These values indicate $4.6\sigma$, $5.3\sigma$, and $5.3\sigma$ tension with respect to the SH0ES measurement, respectively. Several of these results further intensify the existing tension of $4.8\sigma$ between the Planck 2018 results Aghanim et al. (2020) ($H_{0}=67.27\pm 0.60\ \mathrm{km\ s^{-1}Mpc^{-1}}$) and those from SH0ES measurements. The Hubble tension has been extensively investigated in recent years. For comprehensive reviewson this topic, we refer readers to works by Verde et al. (2019); Di Valentino et al. (2021); Abdalla et al. (2022). Given that there is a preference for dynamical dark energy indicated by DESI DR2, modified cosmological models that favor higher values of $H_{0}$ may be incompatible with the late-time dynamical dark energy. This potential contradiction could weaken the statistical preference for such dynamical dark energy scenarios.
@@ II Evidence for dynamical dark energy in the $w_{0}w_{a}$+EDE model
$\log(10^{10}A_{\mathrm{s}})$ $3.069(3.066)\pm 0.013$ $3.068(3.070)\pm 0.013$ 3.069($3.069(3.062)^{+0.012}_{-0.013}$ $3.066(3.071)\pm 0.013$
$n_{\mathrm{s}}$ $0.9870(0.9895)\pm 0.0063$ $0.9867(0.9898)\pm 0.0058$ $0.9870(0.9870)\pm 0.0057$ $0.9853(0.9863)\pm 0.0061$
$\Omega_{\mathrm{b}}h^{2}$ $0.02257(0.02250)\pm 0.00018$ 0.02257($0.02257(0.02240)^{+0.00017}_{-0.00019}$ 0.02256($0.02256(0.02262)^{+0.00016}_{-0.00019}$ $0.02255(0.02249)\pm 0.00019$
$\Omega_{\mathrm{c}}h^{2}$ $0.1282(0.1316)\pm 0.0031$ $0.1286(0.1323)\pm 0.0028$ $0.1284(0.1296)\pm 0.0028$ $0.1284(0.1286)\pm 0.0030$
$\tau_{\mathrm{reio}}$ $0.0602(0.0540)\pm 0.0070$ $0.0595(0.0572)\pm 0.0071$ 0.0604($0.0604(0.0568)^{+0.0063}_{-0.0074}$ 0.0583($0.0583(0.0606)^{+0.0063}_{-0.0070}$
$H_{0}[\mathrm{km\ s^{-1}Mpc^{-1}}]$ 71.44($71.44(72.1)^{+0.87}_{-0.74}$ $71.38(72.32)\pm 0.78$ $71.36(71.43)\pm 0.73$ 71.14($71.14(71.15)^{+0.83}_{-0.75}$
$f_{\mathrm{EDE}}$ 0.104($0.104(0.129)^{+0.027}_{-0.022}$ 0.105($0.105(0.133)^{+0.024}_{-0.021}$ 0.105($0.105(0.112)^{+0.023}_{-0.021}$ 0.100($0.100(0.106)^{+0.027}_{-0.023}$
$\log_{10}a_{c}$$-3.63(-3.59)^{+0.12}_{-0.081}$$-3.62(-3.58)^{+0.12}_{-0.089}$$-3.64(-3.65)^{+0.12}_{-0.072}$$-3.61(-3.73)^{+0.14}_{-0.081}$
$\chi^{2}_{\mathrm{bestfit}}$ 11004.10 12412.27 11034.99 12655.33
@@
$\log(10^{10}A_{\mathrm{s}})$ $3.054(3.052)\pm 0.015$ $3.057(3.069)\pm 0.014$ $3.056(3.062)\pm 0.014$ $3.059(3.047)\pm 0.014$
$n_{\mathrm{s}}$ $0.9772(0.9769)\pm 0.0072$ $0.9827(0.9898)\pm 0.0067$ $0.9832(0.9840)\pm 0.0073$ $0.9849(0.9851)\pm 0.0073$
$\Omega_{\mathrm{b}}h^{2}$ $0.02243(0.02239)\pm 0.00018$ $0.02251(0.02264)\pm 0.00019$ $0.02252(0.02239)\pm 0.00019$ $0.02255(0.02250)\pm 0.00020$
$\Omega_{\mathrm{c}}h^{2}$ $0.1254(0.1266)\pm 0.0034$ $0.1284(0.1314)\pm 0.0028$ $0.1288(0.1284)\pm 0.0031$ $0.1296(0.1300)\pm 0.0031$
$\tau_{\mathrm{reio}}$ 0.0552($0.0552(0.0526)^{+0.0069}_{-0.0080}$ 0.0545($0.0545(0.0557)^{+0.0067}_{-0.0077}$ 0.0537($0.0537(0.0560)^{+0.0062}_{-0.0074}$ 0.0546($0.0546(0.0477)^{+0.0065}_{-0.0075}$
$H_{0}[\mathrm{km\ s^{-1}Mpc^{-1}}]$ $71.60(72.10)\pm 1.1$ $70.78(71.83)\pm 0.72$ $70.44(70.38)\pm 0.84$ $70.37(70.44)\pm 0.85$
$f_{\mathrm{EDE}}$ $0.065(0.076)\pm 0.032$ 0.093($0.093(0.126)^{+0.029}_{-0.025}$ 0.096($0.096(0.095)^{+0.031}_{-0.025}$ 0.104($0.104(0.108)^{+0.029}_{-0.025}$
$\log_{10}a_{c}$$-3.62(-3.63)^{+0.20}_{-0.10}$$-3.59(-3.59)^{+0.15}_{-0.072}$$-3.61(-3.59)^{+0.15}_{-0.086}$$-3.63(-3.58)^{+0.16}_{-0.078}$
$w_{0}$$-1.020(-1.037)\pm 0.10$$-0.896(-0.933)\pm 0.053$$-0.841(-0.847)\pm 0.072$$-0.823(-0.808)\pm 0.053$
$w_{a}$$-0.16(-0.12)^{+0.31}_{-0.27}$$-0.41(-0.22)^{+0.22}_{-0.19}$$-0.57(-0.54)\pm 0.24$$-0.61(-0.63)\pm 0.21$
$\chi^{2}_{\mathrm{bestfit}}$ 11004.00 12407.17 11029.85 12645.23
@@
$\Delta\chi^{2}_{\mathrm{MAP}}$$-1.17$$-5.09$$-5.13$$-11.45$
Significance $0.6\sigma$ $1.8\sigma$ $1.8\sigma$ $2.9\sigma$
@@
    We impose a Gaussian prior of $H_{0}=73.04\pm 1.04$ $\mathrm{km\ s^{-1}Mpc^{-1}}$, which is reported in Ref. Riess et al. (2022).
@@ III Discussion
As shown in 3, the joint analysis using CMB+DESI DR2 BAO+SH0ES yields dark energy constraints consistent with a cosmological constant within $1\sigma$ confidence level. Furthermore, when incorporating Pantheon Plus, Union3, or DESY5 supernova samples alongside CMB+DESI DR2 BAO+SH0ES, we observe preferences for dynamical dark energy at $1.8\sigma$, $1.8\sigma$ and $2.9\sigma$, respectively. Compared to the constraints on $w_{0}w_{a}$ model derived from CMB+DESI DR2 BAO+Pantheon Plus/Union3/DESY5 data (which indicate a preference for dynamical dark energy at $2.8\sigma$, $3.8\sigma$ and $4.2\sigma$, respectively), the statistical preference for dynamical dark energy is significantly reduced.

### v2

@@ Abstract
Recent findings from the Dark Energy Spectroscopic Instrument (DESI) Data Release 2 (DR2) favor a dynamical dark energy characterized by a phantom crossing feature. This result also implies a lower value of the Hubble constant, thereby intensifying the so-called Hubble tension. To alleviate the Hubble tension, we consider the early dark energy and explore its impact on the evidence for dynamical dark energy including the Hubble constant calibrated by the SH0ES collaboration. We find that incorporating SH0ES prior with CMB, DESI DR2 BAO and Pantheon Plus/Union3/DESY5 data reduces the preference to dynamical dark energy to $1.5\sigma/1.4\sigma/2.4\sigma$ level, respectively. Our results suggest a potential tension between the Hubble constant $H_{0}$ of the SH0ES measurement and the phantom-to-quintessence transition in dark energy favored by DESI DR2 BAO data.
@@ I Introduction
Furthermore, the constrains on the $w_{0}w_{a}$ model indicate an evolutionary behavior in the dark energy equation of state, transitioning from $w<-1$ (phantom-like Caldwell (2002)) at high redshifts to $w>-1$ (quintessence-like Ratra and Peebles (1988); Ferreira and Joyce (1998); Wetterich (1988); Scherrer and Sen (2008)) at low redshifts. This behavior implies that the dark energy density decreases with the expansion of the Universe at low redshifts, ultimately resulting in a lower value for the Hubble constant $H_{0}$. As reported in Abdul Karim et al. (2025b), the Hubble constant is determined to be $H_{0}=63.6^{+1.6}_{-2.1}\;\mathrm{km\ s^{-1}Mpc^{-1}}$ in the $w_{0}w_{a}$ model, constrained by CMB+DESI DR2 BAO. This value exhibits a tension of $5.3\sigma$ with the SH0ES measurement ($H_{0}=73.17\pm 0.86$ $\mathrm{km\ s^{-1}Mpc^{-1}}$) Breuval et al. (2024). Furthermore, when considering data combinations such as CMB+DESI DR2 BAO+Pantheon Plus/Union3/DESY5, the corresponding constraints are found to be $H_{0}=67.51\pm 0.59$, $65.91\pm 0.84$, and $66.74\pm 0.56\;\mathrm{km\ s^{-1}Mpc^{-1}}$. These values indicate $5.4\sigma$, $6.0\sigma$, and $6.3\sigma$ tension with respect to the SH0ES measurement, respectively. Several of these results further intensify the existing tension of $5.6\sigma$ between the Planck 2018 results Aghanim et al. (2020) ($H_{0}=67.27\pm 0.60\ \mathrm{km\ s^{-1}Mpc^{-1}}$) and those from SH0ES measurements. The Hubble tension has been extensively investigated in recent years. For comprehensive reviews on this topic, we refer readers to works by Verde et al. (2019); Di Valentino et al. (2021); Abdalla et al. (2022). Given that there is a preference for dynamical dark energy indicated by DESI DR2, modified cosmological models that favor higher values of $H_{0}$ may be incompatible with the late-time dynamical dark energy. This potential contradiction could weaken the statistical preference for such dynamical dark energy scenarios.
@@ II Evidence for dynamical dark energy in the $w_{0}w_{a}$+EDE model
$\log(10^{10}A_{\mathrm{s}})$ $3.071(3.072)\pm 0.013$ $3.069(3.067)\pm 0.013$ $3.069(3.071)\pm 0.013$ $3.068(3.066)\pm 0.013$
$n_{\mathrm{s}}$ $0.9891(0.9923)\pm 0.0060$ $0.9885(0.9889)\pm 0.0061$ $0.9888(0.9848)\pm 0.0057$ $0.9881(0.9890)\pm 0.0057$
$\Omega_{\mathrm{b}}h^{2}$ $0.02261(0.02253)\pm 0.00019$ $0.02259(0.02248)\pm 0.00020$ $0.02259(0.02246)\pm 0.00019$ $0.02258(0.02256)\pm 0.00019$
$\Omega_{\mathrm{c}}h^{2}$ $0.1293(0.1330)\pm 0.0028$ $0.1297(0.1310)\pm 0.0029$ $0.1297(0.1295)\pm 0.0027$ $0.1299(0.1303)\pm 0.0028$
$\tau_{\mathrm{reio}}$ 0.0603($0.0603(0.0578)^{+0.0066}_{-0.0074}$ $0.0591(0.0563)\pm 0.0071$ $0.0593(0.0575)\pm 0.0070$ $0.0584(0.0568)\pm 0.0070$
$H_{0}[\mathrm{km\ s^{-1}Mpc^{-1}}]$ $71.82(72.45)\pm 0.74$ $71.74(72.02)\pm 0.77$ $71.77(71.77)\pm 0.69$ $71.65(71.86)\pm 0.72$
$f_{\mathrm{EDE}}$ 0.115($0.115(0.139)^{+0.022}_{-0.019}$ 0.115($0.115(0.124)^{+0.023}_{-0.019}$ 0.115($0.115(0.112)^{+0.022}_{-0.018}$ 0.115($0.115(0.120)^{+0.022}_{-0.020}$
$\log_{10}a_{c}$$-3.63(-3.58)^{+0.12}_{-0.072}$$-3.62(-3.57)^{+0.12}_{-0.072}$$-3.62(-3.56)^{+0.11}_{-0.067}$$-3.62(-3.61)^{+0.13}_{-0.066}$
$\chi^{2}_{\mathrm{bestfit}}$ 11005.10 12411.71 11033.18 12656.20
@@
$\log(10^{10}A_{\mathrm{s}})$ $3.055(3.059)\pm 0.014$ $3.060(3.055)\pm 0.014$ $3.060(3.054)\pm 0.014$ $3.060(3.067)\pm 0.014$
$n_{\mathrm{s}}$ 0.9784($0.9784(0.9845)^{+0.0072}_{-0.0083}$ $0.9857(0.9841)\pm 0.0070$ $0.9863(0.9846)\pm 0.0071$ $0.9875(0.9907)\pm 0.0069$
$\Omega_{\mathrm{b}}h^{2}$ 0.02247($0.02247(0.02237)^{+0.00018}_{-0.00020}$ $0.02256(0.02245)\pm 0.00019$ $0.02257(0.02236)\pm 0.00020$ $0.02258(0.02250)\pm 0.00020$
$\Omega_{\mathrm{c}}h^{2}$ $0.1259(0.1292)\pm 0.0033$ 0.1300($0.1300(0.1318)^{+0.0027}_{-0.0031}$ $0.1307(0.1325)\pm 0.0030$ $0.1314(0.1344)\pm 0.0030$
$\tau_{\mathrm{reio}}$ $0.0553(0.0548)\pm 0.0072$ $0.0548(0.0520)\pm 0.0072$ $0.0542(0.0483)\pm 0.0072$ $0.0539(0.0539)\pm 0.0071$
$H_{0}\mathrm{km\ s^{-1}Mpc^{-1}}$ $72.16(72.57)\pm 0.83$ $71.40(71.31)\pm 0.70$ $71.23(71.25)\pm 0.74$ $71.01(71.28)\pm 0.73$
$f_{\mathrm{EDE}}$ $0.070(0.102)\pm 0.031$ $0.109(0.119)\pm 0.026$ 0.114($0.114(0.126)^{+0.028}_{-0.025}$ 0.120($0.120(0.142)^{+0.027}_{-0.024}$
$\log_{10}a_{c}$$-3.61(-3.59)^{+0.20}_{-0.10}$$-3.61(-3.59)^{+0.15}_{-0.063}$$-3.61(-3.55)^{+0.14}_{-0.052}$$-3.60(-3.59)^{+0.13}_{-0.051}$
$w_{0}$$-1.057(-1.03)\pm 0.094$$-0.908(-0.878)\pm 0.048$$-0.869(-0.844)\pm 0.069$$-0.833(-0.817)\pm 0.052$
$w_{a}$$-0.07(-0.081)^{+0.28}_{-0.25}$$-0.38(-0.47)^{+0.20}_{-0.18}$$-0.49(-0.54)\pm 0.24$0.57(−$-0.57(-0.58)^{+0.22}_{-0.20}$
$\chi^{2}_{\mathrm{bestfit}}$ 11003.21 12408.87 11030.17 12646.80
@@
$\Delta\chi^{2}_{\mathrm{MAP}}$$-3.18$$-3.91$$-3.80$$-8.31$
Significance $1.3\sigma$ $1.5\sigma$ $1.4\sigma$ $2.4\sigma$
@@
    We impose a Gaussian prior of $H_{0}=73.17\pm 0.86$ $\mathrm{km\ s^{-1}Mpc^{-1}}$, which is reported in Ref. Breuval et al. (2024).
@@ III Discussion
As shown in Table 3, the joint analysis using CMB+DESI DR2 BAO+SH0ES yields a preference for late-time dynamical dark energy at $1.3\sigma$ confidence level. Furthermore, when incorporating Pantheon Plus, Union3, or DESY5 supernova samples alongside CMB+DESI DR2 BAO+SH0ES, we observe preferences for dynamical dark energy at $1.5\sigma$, $1.4\sigma$ and $2.4\sigma$, respectively. Compared to the constraints on $w_{0}w_{a}$ model derived from CMB+DESI DR2 BAO+Pantheon Plus/Union3/DESY5 data (which indicate a preference for dynamical dark energy at $2.8\sigma$, $3.8\sigma$ and $4.2\sigma$, respectively), the statistical preference for dynamical dark energy is significantly reduced.